30x^3 - 20x^2 = 10x
Answer by checkley77(12844)

30x^3 - 20x^2 = 10x
30x^3-20x^2-10x=0
10x(3x^2-2x-1)=0
10x=0
x=0 ans
(3x+1)(x-1)=0
3x+1=0
3x=-1
x=-1/3 ans.
x-1=0
x=1 ans.